# Thermodynamically consistent Cahn-Hilliard-Navier-Stokes equations using the metriplectic dynamics formalism

### v2

@@ 2.4.2 Fluxes and affinities for the GNS system
Comparison of (77) and (83) suggests we require the affinities associated with $\mathbf{m}$, $e$, and $\tilde{c}$. The conventional choices for these affinities are $\nabla(1/T)$, $\nabla(-\mathbf{v}/T)$, and $\nabla(-\mu/T)$, respectively (de Groot and Mazur, 1962). However, examination of (10) or (66) suggests using instead $\nabla T$, $\nabla\mathbf{v}$, and $\nabla\mu$, as was done in Coquinot and Morrison (2020).
@@ 3.1 Hamiltonian and Entropy functional forms
The phenomenon of material transport along an interface is known as the Marangoni effect. The presence of a surface tension gradient naturally induces the migration of particles, moving from regions of low tension to those of high tension. This gradient can be triggered by a concentration gradient (or also a temperature gradient). In two-phase theory the interface between phases is regarded as being diffuse. According to the work of Taylor and Cahn (1998), one can model the diffuse interface by a single order parameter, say $\phi$ and with a free energy functional,
@@
In the isotropic surface energy case Guo and Lin (2015) develop a phase-field model for two-phase flow, which is thermodynamically consistent. The modeling based on a non-classical choice of energy and entropy, given respectively by
@@
Alternatively, Anderson et al. (2000) propose a model of phase-field of solidification with convection, the model permits the interface to have an anisotropic surface energy. The choice of energy and entropy are given by
@@
where $u$, $s$ and $f$ stand for the classical specific internal energy, entropy, and free energy, respectively, the coefficients $\lambda_{s}$ and $\lambda_{u}$ are constant parameters, and $\lambda_{f}(T)$ is a parameter depending on the temperature that will lead to anisotropic surface energy effects. He have defined the total densities $e^{a}_{\mathrm{Total}}$ and $\sigma^{a}_{\mathrm{Total}}$ for later use. The parameter $a$ takes on two values: $a=0$ reduces (97) and (98) to the expressions of Anderson et al. (2000), where we set $\epsilon^{2}_{E}=\lambda_{u}$ and $\epsilon^{2}_{S}=-\lambda_{s}$, while for $a=1$ they reduce to those used by Guo and Lin (2015) provided the choice of an isotropic surface energy is assumed, viz., $\Gamma(\nabla c)=|\nabla c|$. Thus, as is clear from (98), (97), and (99) that the dimensions of $\lambda_{f}$, $\lambda_{s}$, and $\lambda_{u}$ are either specific or volumetric depending on the case. As usual, we have the thermodynamic relation
@@
To summarize, our expressions (97) and (98) generalize the model studied by Guo and Lin (2015) by including $\Gamma$, which accounts for anisotropic surface energy effects, while our expressions generalize the model of Anderson et al. (2000) by including the factors of $\rho$ in the integrands making all quantities in the integrands specific quantities multiplied by the density, giving rise to more general sources of energy.
@@
Upon setting $a=1$ (the case of Guo and Lin, 2015), this reduces to
@@
an expression that differs from that in Guo and Lin (2015) unless $\lambda_{f}\rho$ is constant. If this is the case and we choose a classical $\mu=c^{3}-c$, corresponding to the quartic Laudau potential, we obtain
@@
the chemical potential of Cahn and Hilliard who indeed make these assumptions (cf. page 267 of Cahn and Hilliard, 1958). For $\alpha=0$ (the case of Anderson et al., 2000), we have
@@ 3.2 Noncanonical Poisson bracket of the Cahn-Hilliard-Euler system
Let us now comment on these two Hamiltonian systems. By construction both the $a=1$ and $a=0$ systems conserve their Hamiltonians and entropies, as given by (98) and (97) with $a=1$ and $a=0$ , respectively. Both systems have momentum equations containing a term describing anisotropic surface energy (capillary) effects. The $a=0$ system of (129)–(132) is identical to the ideal limit of that given in the work of Anderson et al. (2000). Upon choosing $\Gamma(\nabla c)=|\nabla c|$, the $a=1$ system of (124)–(127) should correspond to the ideal limit of that of Guo and Lin (2015), but it does not. In fact the system of Guo and Lin (2015) in this limit does not conserve energy. Moreover, the capillary effect in their momentum equation (equation (3.40)), which should be replaced by (124) with $\Gamma(\nabla c)=|\nabla c|$, vanishes in the one-dimensional limit. Since such surface effects are determined by mean or weighted mean curvature (Taylor, 1992), it is clear that this is physically untenable. Fortunately, our method provides a simple fix to their equations, while showing how to generalize them to include anisotropic surface effects.
@@ 4 Summary and Conclusions
From the examples presented, it is clear that the 4-bracket formalism can be applied to obtain a wide variety of dynamical systems in various fields. In fact it was recently applied to obtain generalized collision operators in kinetic theory (Sato and Morrison, 2024) and a thermodynamically consistent model for radiation hydrodynamics (Tran et al., 2024).
The metriplectic 4-bracket formalism also provides an avenue for designing structure preserving numerical algorithms (see e.g. Morrison, 2017). Any discretization that preserves the symmetries of the 4-bracket, which is not a difficult task, will be thermodynamically consistent on the semi-discrete level, i.e. produce a set of ordinary differential equations that conserve energy and produce entropy.
@@ Appendix A Semidefinite curvature
Finite-dimensional versions of these two lemmas were first reported in Morrison and Updike (2023).

### v3

@@ 2.4.2 Fluxes and affinities for the GNS system
Comparison of (77) and (83) suggests we require the affinities associated with $\mathbf{m}$, $e$, and $\tilde{c}$. The conventional choices for these affinities are $\nabla(1/T)$, $\nabla(-\mathbf{v}/T)$, and $\nabla(-\mu/T)$, respectively [16]. However, examination of (10) or (66) suggests using instead $\nabla T$, $\nabla\mathbf{v}$, and $\nabla\mu$, as was done in Coquinot and Morrison [6].
@@ 3.1 Hamiltonian and Entropy functional forms
The phenomenon of material transport along an interface is known as the Marangoni effect. The presence of a surface tension gradient naturally induces the migration of particles, moving from regions of low tension to those of high tension. This gradient can be triggered by a concentration gradient (or also a temperature gradient). In two-phase theory the interface between phases is regarded as being diffuse. According to the work of Taylor and Cahn [35], one can model the diffuse interface by a single order parameter, say $\phi$ and with a free energy functional,
@@
In the isotropic surface energy case Guo and Lin [17] develop a phase-field model for two-phase flow, which is thermodynamically consistent. The modeling based on a non-classical choice of energy and entropy, given respectively by
@@
Alternatively, Anderson et al. [2] propose a model of phase-field of solidification with convection, the model permits the interface to have an anisotropic surface energy. The choice of energy and entropy are given by
@@
where $u$, $s$ and $f$ stand for the classical specific internal energy, entropy, and free energy, respectively, the coefficients $\lambda_{s}$ and $\lambda_{u}$ are constant parameters, and $\lambda_{f}(T)$ is a parameter depending on the temperature that will lead to anisotropic surface energy effects. He have defined the total densities $e^{a}_{\mathrm{Total}}$ and $\sigma^{a}_{\mathrm{Total}}$ for later use. The parameter $a$ takes on two values: $a=0$ reduces (97) and (98) to the expressions of Anderson et al. [2], where we set $\epsilon^{2}_{E}=\lambda_{u}$ and $\epsilon^{2}_{S}=-\lambda_{s}$, while for $a=1$ they reduce to those used by Guo and Lin [17] provided the choice of an isotropic surface energy is assumed, viz., $\Gamma(\nabla c)=|\nabla c|$. Thus, as is clear from (98), (97), and (99) that the dimensions of $\lambda_{f}$, $\lambda_{s}$, and $\lambda_{u}$ are either specific or volumetric depending on the case. As usual, we have the thermodynamic relation
@@
To summarize, our expressions (97) and (98) generalize the model studied by Guo and Lin [17] by including $\Gamma$, which accounts for anisotropic surface energy effects, while our expressions generalize the model of Anderson et al. [2] by including the factors of $\rho$ in the integrands making all quantities in the integrands specific quantities multiplied by the density, giving rise to more general sources of energy.
@@
Upon setting $a=1$ [the case of 17], this reduces to
@@
an expression that differs from that in Guo and Lin [17] unless $\lambda_{f}\rho$ is constant. If this is the case and we choose a classical $\mu=c^{3}-c$, corresponding to the quartic Laudau potential, we obtain
@@
the chemical potential of Cahn and Hilliard who indeed make these assumptions [cf. page 267 of 4]. For $\alpha=0$ [the case of 2], we have
@@ 3.2 Noncanonical Poisson bracket of the Cahn-Hilliard-Euler system
where recall from (103), $\xi=\partial\Gamma/\partial\mathbf{p}$
Let us now comment on these two Hamiltonian systems. By construction both the $a=1$ and $a=0$ systems conserve their Hamiltonians and entropies, as given by (98) and (97) with $a=1$ and $a=0$ , respectively. Both systems have momentum equations containing a term describing anisotropic surface energy (capillary) effects. The $a=0$ system of (129)–(132) is identical to the ideal limit of that given in the work of Anderson et al. [2]. Upon choosing $\Gamma(\nabla c)=|\nabla c|$, the $a=1$ system of (124)–(127) should correspond to the ideal limit of that of Guo and Lin [17], but it does not. In fact the system of Guo and Lin [17] in this limit does not conserve energy. Moreover, the capillary effect in their momentum equation (equation (3.40)), which should be replaced by (124) with $\Gamma(\nabla c)=|\nabla c|$, vanishes in the one-dimensional limit. Since such surface effects are determined by mean or weighted mean curvature [34], it is clear that this is physically untenable. Fortunately, our method provides a simple fix to their equations, while showing how to generalize them to include anisotropic surface effects.
@@ 4 Summary and Conclusions
From the examples presented, it is clear that the 4-bracket formalism can be applied to obtain a wide variety of dynamical systems in various fields. In fact it was recently applied to obtain generalized collision operators in kinetic theory [32] and a thermodynamically consistent model for radiation hydrodynamics [37]. Although incompressible flows don’t have the usual thermodynamics associated with compression and pressure, they can be included in the metriplectic formalism by using the techniques of Chandre et al. (2013).
The metriplectic 4-bracket formalism also provides an avenue for designing structure preserving numerical algorithms [see e.g. 27]. Any discretization that preserves the symmetries of the 4-bracket, which is not a difficult task, will be thermodynamically consistent on the semi-discrete level, i.e. produce a set of ordinary differential equations that conserve energy and produce entropy.
@@ Appendix A Semidefinite curvature
Finite-dimensional versions of these two lemmas were first reported in Morrison and Updike [30].